In 2019, X Company sold 5,150 units of its only product. Total revenue was $1,351,360, total variable costs were $567,530, and total fixed costs were $666,000. In 2020, there are only two expected changes: variable costs will decrease by $4.60 per unit, and fixed costs will decrease by $42,000. How many units must X Company sell in 2020 in order to earn $30,000?
Answer #1

Sales price per unit was=(1,351,360/5150)=$262.4 per unit

Variable costs per unit was=(567,530/5150)=$110.2 per unit

New variable costs=(110.2-4.6)=$105.6 per unit

New fixed cost=(666,000-42000)=$624,000

Contribution margin=Sales-Variable cost  

=(262.4-105.6)=$156.8 per unit

Total Contribution margin required=Fixed cost+Target profit

=624,000+30,000=654000

Hence units to be sold=654000/156.8

=4170.92 units(Approx).